A bag containing 5 red balls, 3 blue balls and 2 yellow balls. A ball is drawn and not replaced. A second ball is drawn. Find the probabilty of drawing:
a) Two red balls
b) One blue ball and one yellow ball
c) Two yellow balls
d) A red ball and a blue ball

Answer by jim_thompson5910(35256)

a)

P(Two red) = P(Red)*P(Red after one is drawn) = (5/10)(4/9) = 20/180 = 1/9

b)

P(One blue and one yellow) = P(blue) * P(yellow after one draw) = (3/10)(2/9) = 6/90 = 1/15

c)

P(Two yellow) = P(yellow)*P(yellow after one is drawn) = (2/10)(1/9) = 2/90 = 1/45


d)

P(One red and one blue) = P(red) * P(blue after one draw) = (5/10)(3/9) = 15/90 = 1/6